chocolateluva77:

can smo explain how to put this quadratic function into standard form? pls explain how u got the steps to the solution ty

chocolateluva77:

snowflake0531:

to get a function into standard form, you need to get it to ax^2 + bx + c I suggest simplifying (x-1)^2 using the (a-b)^2 = a^2 - 2ab + b^2 formula, then adding 1 to it

chocolateluva77:

@snowflake0531 where do u add 1 to?

chocolateluva77:

@snowflake0531 is it (x-1)^2 = x^2-2(1)(-1)+(-1)^2+1?

snowflake0531:

\[(x-1)^2+1=x^2-2x+1+1=x^2-2x+2\]

chocolateluva77:

@snowflake0531 ty!

snowflake0531:

yw